# On the Equivalence of Koopman Eigenfunctions and Commuting Symmetries

### v3

@@ Abstract
The Koopman operator framework offers a way to represent a nonlinear system as a linear one. The key to this simplification lies in the identification of eigenfunctions. While various data-driven algorithms have been developed for this problem, a theoretical characterization of Koopman eigenfunctions from geometric properties of the flow is still missing. This paper provides such a characterization by establishing an equivalence between a set of Koopman eigenfunctions and a set of commuting symmetries—both assumed to span the tangent spaces at every point on a simply connected open set. Based on this equivalence, we build an explicit and convergent formula for the principal Koopman eigenfunctions defined on the region of attraction of a locally asymptotically stable equilibrium point, thereby offering a constructive formula to compute Koopman eigenfunctions.
@@ I Introduction
Nonlinear dynamical systems are often studied through the evolution of functions of the state, i.e., the so-called observables. The Koopman operator framework offers a linear perspective on nonlinear dynamics. It describes the evolution of observables via an infinite-dimensional linear operator. Koopman eigenfunctions, in particular, serve as powerful tools for spectral analysis, dimensionality reduction, and model discovery [1, 2].
Despite their importance, the computation of Koopman eigenfunctions remains challenging. Data-driven methods, such as extended dynamic mode decomposition (EDMD) and related variants [3, 4, 5], provide practical approximations from simulation or experimental data. Recent contributions have introduced finite-data error bounds, which quantify the reliability of these approximations from both a learning-theoretic perspective [6] and control-oriented contexts [7]. These results underscore the potential of Koopman-based methods. At the same time, they also highlight a fundamental limitation arising from the absence of analytical characterizations of eigenfunctions.
One emerging direction to address this limitation is to incorporate system symmetries into Koopman learning. By enforcing equivariance or symmetry constraints, researchers have extended Koopman-based models to dynamical systems with geometric or physical structure [8, 9, 10, 11, 12, 13, 14, 15]. These approaches demonstrate that symmetry constraints can improve the stability of data-driven algorithms and enhance their interpretability. Nevertheless, most existing methods impose symmetry constraints algorithmically, without establishing a principled connection between Koopman eigenfunctions and intrinsic system symmetries.
@@
This paper establishes a connection between the two perspectives by establishing a direct equivalence between Koopman eigenfunctions and commuting symmetries. Specifically, we show that on a simply connected open domain, the gradients of Koopman eigenfunctions are mapped bijectively to commuting symmetry generators that span the tangent spaces. This equivalence provides a structural characterization of Koopman eigenfunctions that complements existing data-driven methods. Based on it, we also derive an explicit formula for Koopman eigenfunctions. We then prove that it converges uniformly on compact subsets of a region of attraction. In summary, the main contributions of this paper are:
@@
This paper is organized as follows. Section II covers geometric preliminaries. Section III connects Koopman eigenfunctions to commuting symmetries. Section IV derives the formula for Koopman eigenfunctions. Section V offers concluding remarks.
@@
The imaginary unit is $j$. A function is denoted by lower-case italic (e.g., $\psi$). A vector field [18] is denoted by upper-case italic (e.g., $F$). The exceptions are $U$, AA, and $K$, which are subsets of $\mathbb{R}^{n}$. The zero vector in $\mathbb{R}^{n}$ is $0_{n}$. The standard basis of $\mathbb{R}^{n}$ is $(\mathbf{e}_{1},\ldots,\,\mathbf{e}_{n})$. $(\cdot)^{*}$ denotes conjugate transpose, and $(\cdot)^{\mathsf{T}}$ denotes transpose. The gradient of a complex-valued function $g\mathrel{\mathop{\ordinarycolon}}\mathbb{R}^{n}\to\mathbb{C}$ is a column vector such that
$Dg(\mathbf{x})\mathbf{v}=\nabla g(\mathbf{x})^{*}\mathbf{v},\quad(\mathbf{v})\in\mathbb{R}^{n}.$
The Jacobian of a vector field $X$ on $\mathbb{R}^{n}$ is a square matrix such that
$DX(\mathbf{x})\mathbf{v}=\nabla X(\mathbf{x})\mathbf{v},\quad\mathbf{v}\in\mathbb{R}^{n}.$
@@ II Preliminaries
$\dot{\mathbf{x}}=F(\mathbf{x}),\,\mathbf{x}\in\mathbb{R}^{n},$ (1)
starting from the initial condition $\mathbf{x}_{0}\in\mathbb{R}^{n}$. If $F(\mathbf{x})$ is $\mathcal{C}^{\infty}$ smooth, then the trajectory from every initial condition is unique. The flow $\Phi_{F}(t,\mathbf{x})$ is said to be complete if, for each $\mathbf{x}\in\mathbb{R}^{n}$, it belongs to $\mathbb{R}^{n}$ for all $t\in\mathbb{R}$; that is, it is invariant in $\mathbb{R}^{n}$. For simplicity, we assume that the flow $\Phi_{F}(t,\mathbf{x})$ is complete; this holds for many physically relevant models [19]. We will also use the notation $\Phi_{F}^{t}(\mathbf{x})$ for $\Phi_{F}(t,\mathbf{x})$.
@@
The Lie derivative $\mathcal{L}_{F}$ gives the infinitesimal change of a function $\varphi(\mathbf{x})$ along the flow $\Phi_{F}$, independent of coordinate choice. [^1] The coordinates for (1) are the state vector $\mathbf{x}$, according to which the partial derivatives and the gradient are defined. Later we will see that the Lie derivative is the generator of the Koopman operator.
@@ II-A Definition of Trajectory Symmetries
A one-parameter group is a mapping $\mathcal{T}(s)$ from the additive group $\mathbb{R}$ to a Lie group $\mathbb{G}$ that satisfies the group homomorphism, i.e., $\mathcal{T}(r)\mathcal{T}(s)=\mathcal{T}(r+s)$. The infinitesimal generator of $\mathcal{T}(s)$ is $G=\frac{\partial\mathcal{T}}{\partial s}(0)$.
@@
where the group property is used in the first equality. Noting that $\mathcal{T}(0)=\mathrm{Id}$ and (3), we obtain $\mathcal{T}(s)=\mathrm{exp}(sG)$ [18]. The primary example we have in mind is the complete flow $\Phi_{F}^{t}$ of (1), which is a one-parameter group of diffeomorphisms of $\mathbb{R}^{n}$ generated by $F(\mathbf{x})$. Formally, symmetries are defined using local one-parameter groups. These are groups homomorphic to $\mathbb{R}$ only near 0, which accounts for the fact that the flow of a vector field may be incomplete.
@@
$\nabla\Phi_{G}^{s}(\mathbf{x})F(\mathbf{x})=F(\Phi_{G}^{s}(\mathbf{x})),\,\forall\mathbf{x}\in U,\,s\in\mathrm{I}_{\mathbf{x}};$ (4)
@@
$[G,F](\mathbf{x})=\nabla G(\mathbf{x})F(\mathbf{x})-\nabla F(\mathbf{x})G(\mathbf{x})=0_{n},\,\forall\mathbf{x}\in U.$ (5)
@@
By the flow-box theorem [20], near each non-equilibrium point $\mathbf{x}^{\prime}$ of (1), there is a local diffeomorphism $\widehat{\mathbf{z}}=Z^{-1}(\mathbf{x})$ such that, in $\widehat{\mathbf{z}}$ coordinates, the flow is the constant flow $\dot{\widehat{\mathbf{z}}}=\mathbf{e}_{1}$. Then, a set of commuting symmetry generators can be taken as $\mathbf{e}_{i}$ in $\widehat{\mathbf{z}}$ coordinates, which correspond to $\nabla Z(\widehat{\mathbf{z}})\mathbf{e}_{i}$ in $\mathbf{x}$ coordinates. However, identifying global symmetries for a general nonlinear system remains challenging [21, 22, 23, 24].
@@ II-B Commuting and Conservative Local Frames
that is full-rank and commuting, i.e., $E_{i},E_{k}](\mathbf{x})=0_{n}$. The goal is to show that this set can be mapped bijectively to a set of $n$ gradients of the $n$ scalar functions. Conceptually, the trajectory lines of the commuting vector fields define a set of curvilinear coordinates. The coordinate values at each point are exactly the values of the associated scalar functions.
@@
$\nabla X(\mathbf{x})=\nabla X(\mathbf{x})^{*},\,\forall\mathbf{x}\in U.$ (7)
@@ III Koopman Eigenfunctions From Commuting Symmetries
Koopman eigenfunctions for (1) are functions that evolve exponentially along the flow, much like eigenvectors for a linear system. As a result, under full rank assumption, they can serve as coordinates that linearize the trajectories of a nonlinear system [25] on nonlocal sets, capturing nonlocal behavior. Hence, identifying the eigenfunctions has become a central objective in modern nonlinear systems research.
We will show that any set of linearly independent (in the sense of a local frame) Koopman eigenfunctions has an associated commuting local frame of symmetries. Specifically, we will sharpen (8) to ensure that $\big{[}X_{1}(\mathbf{x})\;\cdots\;X_{n}(\mathbf{x})\big{]}$ in the LHS is the gradients of the logarithms of Koopman eigenfunctions for (1) if and only if $\big{[}E_{1}(\mathbf{x})\;\cdots\;E_{n}(\mathbf{x})\big{]}$ in the RHS is a commuting local symmetry frame of symmetries for (1).
@@ III-A Koopman Eigenfunctions
$\mathcal{K}(t)\varphi(\mathbf{x})=[\varphi\circ\Phi_{F}](t,\mathbf{x})=\varphi(\Phi_{F}(t,\mathbf{x})),\,t\in\mathbb{R}.$
@@
A function $\psi\mathrel{\mathop{\ordinarycolon}}U\to\mathbb{C}$ is said to be an open Koopman eigenfunction on the open subset $U\subset\mathbb{R}^{n}$ if, for some $\mu\in\mathbb{C}$ (the so-called Koopman eigenvalue), it holds that
$\mathcal{L}_{F}\psi(\mathbf{x})=\nabla\psi(\mathbf{x})^{*}F(\mathbf{x})=\mu\psi(\mathbf{x}),\,\forall\mathbf{x}\in U.$ (9)
@@ III-B Main Result
From properties ii) and iii) in Lemma 3, Koopman eigenfunctions combine with each other through multiplication, which makes it difficult to define linear independence for a set of $n$ Koopman eigenfunctions. We simplify the problem by considering the logarithms of the Koopman eigenfunctions. More precisely, assuming that the Koopman eigenfunctions are nonzero on $U$, we will characterize the gradients of their logarithms, referred to as conservative linearizing vector fields.
@@
Definition 5.
$\nabla X(\mathbf{x})^{*}F(\mathbf{x})+\nabla F(\mathbf{x})^{*}X(\mathbf{x})=0_{n},\,\forall\mathbf{x}\in U.$ (11)
@@
$\nabla m(\mathbf{x})^{*}F(\mathbf{x})=c,\,\forall\mathbf{x}\in U.$ (12)
To clarify the correspondence to Koopman eigenfunctions, let us define $\psi(\mathbf{x})=\exp(m(\mathbf{x}))$, i.e., the exponential of $m(\mathbf{x})$. Then, we have
@@
and it is easy to see that $\psi(\mathbf{x})$ is a Koopman eigenfunction with the eigenvalue given by the constant $c$ in (12). The main result on the equivalence of Koopman eigenfunctions and commuting symmetries is stated below.
@@
All are assumed to be defined on a simply connected open subset $U\subset\mathbb{R}^{n}$. In addition, the Koopman eigenfunctions can be computed by integrating its gradients, which is well-defined on a simply connected subset $U\subset\mathbb{R}^{n}$ by the Poincaré Lemma. By choosing $U$ to be simply connected, we have avoided the problem that the complex logarithm is multi-valued around a singularity, i.e., points where the Koopman eigenfunction is zero. The singularities represent limit sets of the system. If these singularities are allowed to puncture $U$, then the logarithms of Koopman eigenfunctions may become multi-valued around each singularity, whereas the Koopman eigenfunctions are known to be at least $\mathcal{C}^{1}$ around a stable equilibrium point [27].
Remark 3 (Comparison to flow-box).
Remark 1 states that, if there is a commuting local frame of symmetries $\big{[}E_{1}(\mathbf{x})\;\cdots\;E_{n}(\mathbf{x})\big{]}$ for (1), then there exist inverse coordinates $\widehat{\mathbf{z}}=Z^{-1}(\mathbf{x})$ whose gradients $\nabla\widehat{z}_{i}(\mathbf{x})$ coincide with the local frame of commuting symmetries. In comparison, Proposition 1 states that $\nabla\mathbf{z}=\nabla Z(\mathbf{x})$ are conservative linearizing vector fields for (1). The latter correspond to gradients of logarithms of Koopman eigenfunctions, and offer nonlocal representations of (1) as a linear system (at least on the simply connected open set $U$).
@@
Though the flow map is relatively straightforward to construct from n1n-1 commuting symmetries as we have shown above, it is difficult to compute the commuting symmetries from the flow map. This problem involves finding a partial order for the subset $U\subset\mathbb{R}^{n}$ by the flow map and becomes much simpler if $U$ is the region of attraction of a locally asymptotically stable equilibrium point, as we demonstrate next.

### v4

@@ Abstract
The Koopman operator framework offers a way to represent a nonlinear system as a linear one. The key to this simplification lies in the identification of eigenfunctions. While various data-driven algorithms have been developed for this problem, a theoretical characterization of Koopman eigenfunctions from geometric properties of the flow is still missing. This paper provides such a characterization by establishing an equivalence between a set of Koopman eigenfunctions and a set of commuting symmetries—both assumed to span the tangent spaces at every point on a simply connected open set. Based on this equivalence, we derive an explicit formula for the principal Koopman eigenfunctions and prove its uniform convergence on the region of attraction of a locally asymptotically stable equilibrium point, thereby offering a constructive method for computing Koopman eigenfunctions.
@@ I Introduction
Nonlinear dynamical systems are often studied through the evolution of functions of the state, i.e., observables. The Koopman operator framework offers a linear perspective on nonlinear dynamics. It describes the evolution of observables via an infinite-dimensional linear operator. Koopman eigenfunctions, in particular, serve as powerful tools for spectral analysis, dimensionality reduction, and model discovery [1, 2].
Despite their importance, the computation of Koopman eigenfunctions remains challenging. Data-driven methods, such as extended dynamic mode decomposition (EDMD) and related variants [3, 4, 5], provide practical approximations from simulation or experimental data. Recent contributions have introduced finite-data error bounds, which quantify the reliability of these approximations from both a learning-theoretic perspective [6] and control-oriented contexts [7]. These results underscore the potential of Koopman-based methods. At the same time, they highlight a fundamental limitation arising from the absence of analytical characterizations of eigenfunctions.
One emerging direction to address this limitation is to incorporate system symmetries into Koopman learning. By enforcing equivariance or symmetry constraints, researchers have extended Koopman-based models to dynamical systems with geometric or physical structure [8, 9, 10, 11, 12, 13, 14, 15, 16]. These approaches demonstrate that symmetry constraints can improve the stability of data-driven algorithms and enhance their interpretability. Nevertheless, most existing methods impose symmetry constraints algorithmically, without establishing a principled connection between Koopman eigenfunctions and intrinsic system symmetries.
@@
This paper establishes a connection between the two perspectives by giving a direct equivalence between Koopman eigenfunctions and commuting symmetries. Specifically, we show that on a simply connected open domain, the gradients of Koopman eigenfunctions are mapped bijectively to commuting symmetry generators that span the tangent spaces. This equivalence provides a structural characterization of Koopman eigenfunctions that complements existing data-driven methods. Based on it, we also derive an explicit formula for Koopman eigenfunctions. We then prove that it converges uniformly on compact subsets of a region of attraction. In summary, the main contributions of this paper are:
@@
This paper is organized as follows. Section II covers geometric preliminaries. Section III connects Koopman eigenfunctions to commuting symmetries. Section IV specializes to the region of attraction and derives the computable limit (20). Section V offers concluding remarks.
@@
The imaginary unit is $j$. A function is denoted by lower-case italic (e.g., $\psi$). A vector field [19] is denoted by upper-case italic (e.g., $F$). The exceptions are $U$, AA, and $K$, which are subsets of $\mathbb{R}^{n}$. The zero and one vectors in $\mathbb{R}^{n}$ are $0_{n}$ and $1_{n}$, respectively. The standard basis of $\mathbb{R}^{n}$ is $(\mathbf{e}_{1},\ldots,\,\mathbf{e}_{n})$. $(\cdot)^{*}$ denotes conjugate transpose, and $(\cdot)^{\mathsf{T}}$ denotes transpose. For a complex-valued function $g\mathrel{\mathop{\ordinarycolon}}\mathbb{R}^{n}\to\mathbb{C}$, the gradient $\nabla g(\mathbf{x})$ is a column vector such that
$\frac{\partial g(\mathbf{x})}{\partial\mathbf{x}}\mathbf{v}=\nabla g(\mathbf{x})^{*}\mathbf{v},\quad\forall\mathbf{v}\in\mathbb{R}^{n}.$
For a vector field $X(\mathbf{x})$ on $\mathbb{R}^{n}$, the Jacobian $\nabla X(\mathbf{x})$ is a square matrix such that
$\frac{\partial X(\mathbf{x})}{\partial\mathbf{x}}\mathbf{v}=\nabla X(\mathbf{x})\mathbf{v},\quad\forall\mathbf{v}\in\mathbb{R}^{n}.$
@@ II Preliminaries
$\dot{\mathbf{x}}=F(\mathbf{x}),\quad\mathbf{x}\in\mathbb{R}^{n},$ (1)
starting from the initial condition $\mathbf{x}_{0}\in\mathbb{R}^{n}$. If $F(\mathbf{x})$ is $\mathcal{C}^{\infty}$-smooth, then the trajectory from every initial condition is unique. The flow $\Phi_{F}(t,\mathbf{x})$ is said to be complete if, for each $\mathbf{x}\in\mathbb{R}^{n}$, $\Phi_{F}(t,\mathbf{x})$ belongs to $\mathbb{R}^{n}$ for all $t\in\mathbb{R}$; that is, it is invariant in $\mathbb{R}^{n}$. For simplicity, we assume that the flow $\Phi_{F}(t,\mathbf{x})$ is complete; this holds for many physically relevant models [20]. We will also use the notation $\Phi_{F}^{t}(\mathbf{x})$ for $\Phi_{F}(t,\mathbf{x})$.
@@
The linear operator $\mathcal{L}_{F}$ describes the infinitesimal change of a function $\varphi(\mathbf{x})$ along the flow $\Phi_{F}(t,\mathbf{x})$, independently of coordinate choice. [^1] The coordinates for (1) are the state vector $\mathbf{x}$, according to which the partial derivatives and the gradient are defined. Later we will see that $\mathcal{L}_{F}$ generates the Koopman operator.
@@ II-A Definition of Trajectory Symmetries
A one-parameter group is a mapping $\mathcal{T}(s)$ from the additive group $\mathbb{R}$ to a Lie group $\mathbb{G}$ that is a group homomorphism, i.e., $\mathcal{T}(r)\mathcal{T}(s)=\mathcal{T}(r+s)$. The infinitesimal generator of $\mathcal{T}(s)$ is $G=\frac{\partial\mathcal{T}}{\partial s}(0)$.
@@
where the group property is used in the first equality. Noting that $\mathcal{T}(0)=\mathrm{Id}$ and (3), we obtain $\mathcal{T}(s)=\mathrm{exp}(sG)$ [19]. The primary example we have in mind is the complete flow $\Phi_{F}^{t}$ of (1), which is a one-parameter group of diffeomorphisms of $\mathbb{R}^{n}$ generated by $F(\mathbf{x})$. Formally, symmetries are defined using local one-parameter groups. These are groups defined for $s$ in a neighborhood of 0, which accounts for the fact that the flow of a vector field may be incomplete.
@@
$\nabla\Phi_{G}^{s}(\mathbf{x})F(\mathbf{x})=F(\Phi_{G}^{s}(\mathbf{x})),\quad\forall\mathbf{x}\in U,\,s\in\mathrm{I}_{\mathbf{x}};$ (4)
@@
$[G,F](\mathbf{x})=\nabla G(\mathbf{x})F(\mathbf{x})-\nabla F(\mathbf{x})G(\mathbf{x})=0_{n},\quad\forall\mathbf{x}\in U.$ (5)
@@
By the flow-box theorem [21], near each non-equilibrium point $\mathbf{x}^{\prime}$ of (1), there is a local diffeomorphism $\widehat{\mathbf{z}}=Z^{-1}(\mathbf{x})$ such that, in $\widehat{\mathbf{z}}-$coordinates, the flow is the constant flow $\dot{\widehat{\mathbf{z}}}=\mathbf{e}_{1}$. Then, a set of commuting symmetry generators can be taken as $\mathbf{e}_{i}$ in $\widehat{\mathbf{z}}-$coordinates, which correspond to $\nabla Z(\widehat{\mathbf{z}})\mathbf{e}_{i}$ in $\mathbf{x}$-coordinates. However, identifying global symmetries for a general nonlinear system remains challenging [22, 23, 24, 25].
@@ II-B Commuting and Conservative Local Frames
that form a full-rank, commuting set, i.e., $E_{i},E_{k}](\mathbf{x})=0_{n}$. The goal is to show that the $n$ commuting vector fields can be mapped bijectively to the gradients of $n$ scalar functions. Conceptually, the trajectory lines of the commuting vector fields define a set of curvilinear coordinates. The coordinate values at each point are exactly the values of the associated scalar functions.
@@
$\nabla X(\mathbf{x})=\nabla X(\mathbf{x})^{\mathsf{T}},\quad\forall\mathbf{x}\in U.$ (7)
@@ III Koopman Eigenfunctions From Commuting Symmetries
Koopman eigenfunctions for (1) are functions that evolve exponentially along the flow, much like eigenvectors for a linear system. As a result, under a full-rank assumption, they can serve as coordinates that linearize the nonlinear dynamics on a nonlocal set, capturing nonlocal behavior. Hence, identifying the eigenfunctions has become a central objective in modern nonlinear systems research.
We will show that any set of linearly independent (in the sense of a local frame) Koopman eigenfunctions has an associated commuting local frame of symmetries. Specifically, we will sharpen (8) to prove that $\big{[}X_{1}(\mathbf{x})\;\cdots\;X_{n}(\mathbf{x})\big{]}$ in the LHS is a local frame of gradients of logarithms of Koopman eigenfunctions for (1) if and only if $\big{[}E_{1}(\mathbf{x})\;\cdots\;E_{n}(\mathbf{x})\big{]}$ in the RHS is a commuting local symmetry frame for (1).
@@ III-A Koopman Eigenfunctions
$\mathcal{K}(t)\varphi(\mathbf{x})=[\varphi\circ\Phi_{F}](t,\mathbf{x})=\varphi(\Phi_{F}(t,\mathbf{x})),\quad t\in\mathbb{R}.$
@@
A function $\psi\mathrel{\mathop{\ordinarycolon}}U\to\mathbb{C}$ is said to be an open Koopman eigenfunction for (1) on the open subset $U\subset\mathbb{R}^{n}$ if, for some $\mu\in\mathbb{C}$ (the eigenvalue), it holds that
$\mathcal{L}_{F}\psi(\mathbf{x})=\nabla\psi(\mathbf{x})^{*}F(\mathbf{x})=\mu\psi(\mathbf{x}),\quad\forall\mathbf{x}\in U.$ (9)
@@ III-B Main Result
From properties ii) and iii) in Lemma 3, Koopman eigenfunctions combine with each other through multiplication, which makes it difficult to define linear independence for a set of $n$ Koopman eigenfunctions. We simplify the problem by considering the logarithms of the Koopman eigenfunctions. More precisely, assuming that the Koopman eigenfunctions are nonzero on $U$, we will characterize the gradients of their logarithms, which are referred to as conservative linearizing vector fields.
Definition 5.
$\nabla X(\mathbf{x})^{*}F(\mathbf{x})+\nabla F(\mathbf{x})^{*}X(\mathbf{x})=0_{n},\quad\forall\mathbf{x}\in U.$ (11)
@@
$\nabla m(\mathbf{x})^{*}F(\mathbf{x})=c,\quad\forall\mathbf{x}\in U.$ (12)
To connect $m(\mathbf{x})$ to a Koopman eigenfunction, let us consider the function $\psi(\mathbf{x})=\exp(m(\mathbf{x}))$, i.e., the exponential of $m(\mathbf{x})$. Then, we have
@@
and it is readily seen that $\psi(\mathbf{x})$ is a Koopman eigenfunction with the eigenvalue given by the constant $c$ in (12). The main result on the equivalence of Koopman eigenfunctions and commuting symmetries is stated below.
@@
All are assumed to be defined on a simply connected open subset $U\subset\mathbb{R}^{n}$. In addition, the Koopman eigenfunctions can be computed by integrating their gradients, which is well-defined on a simply connected subset $U\subset\mathbb{R}^{n}$ by the Poincaré lemma. By choosing $U$ to be simply connected, we have avoided the problem that the complex logarithm is multi-valued around a singularity, i.e., points where the Koopman eigenfunction is zero. The singularities represent limit sets of the system. If these singularities are allowed to puncture $U$, then on the punctured domain, the complex logarithms of Koopman eigenfunctions can be multi-valued around each singularity, whereas the Koopman eigenfunctions usually remain single-valued [27].
Remark 3 (Comparison to flow-box theorem).
Recall from Remark 1 that, if there is a commuting local frame of symmetries $\big{[}E_{1}(\mathbf{x})\;\cdots\;E_{n}(\mathbf{x})\big{]}$ for (1), then there exist inverse coordinates $\widehat{\mathbf{z}}=Z^{-1}(\mathbf{x})$ whose gradients $\nabla\widehat{\mathbf{z}}=\nabla Z^{-1}(\mathbf{x})$ coincide with the local frame of commuting symmetries. In comparison, Proposition 1 states that $\nabla\mathbf{z}=\nabla Z(\mathbf{x})$ are conservative linearizing vector fields for (1). The latter correspond to the gradients of logarithms of Koopman eigenfunctions, and offer nonlocal representations of (1) as a linear system (at least valid on the simply connected open set $U$).
@@
Though the flow map is relatively straightforward to construct from n1n-1 commuting symmetries as we have shown above, it is difficult to compute the commuting symmetries from the flow map. This problem involves finding a partial order for the subset $U\subset\mathbb{R}^{n}$ by the flow map and becomes much simpler if $U$ is the region of attraction of a locally asymptotically stable equilibrium point, as we show next.